OpenStudy (anonymous):

Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). g(x) = x3 – x2 – 4x + 4

OpenStudy (anonymous):

g(x) = x^3 – x^2 – 4x + 4

OpenStudy (anonymous):

@No.name Can you help me with this problem? It wants me to find the zeros for that function.

OpenStudy (anonymous):

g(x) = x^3 – x^2 – 4x + 4 The zeros are found by factoring the equation. x^2(x-1) - 4(x -1) = 0 (x^2 -4)(x-1) = 0 (x+2)(x-2)(x-1) = 0 x = 2, -2, and 1 The other key features are: g'(x) gives the slope g''(x) gives the concavity g'(x) = 0 will give the critical points g''(x) = 0 will give the point of inflection Copied from http://openstudy.com/study#/updates/526464c3e4b029b030d9ada1

OpenStudy (anonymous):

Oh Okay, Thanks!